CASIC SESSES Therefore, SAAMAAN L UU 0.299x10° Y abs AS , 1LOKN A. 2000 mm² EA = constant E: 200 gra * * 4m + 4m x

analysis of two dimensional trusses (matrix)
Answer #1

Please find the following pictures for the solution to given problem.stiffness matrix Member 6 0 =0 c 1 s 0 = 8m. 30 AE r oes -c-est K: E 1-c-es -- c es on - Les 5. cs s ai Az A3 n on 07: 05 0.1Member- (= 8.94mg 0.89 dc=0.45 K: AE 1 10.79 0.40 -0.49 -0.407 8.94 .0.40 0.20 -0.40 -0.20 -0.79 -0.40 0.79 0.40. 1-0.40 -0.2Global stiffness matrix 10.303 [x] - A6% 0-135 -135 31-0195 4 0 5-0.09 6 -0.09 -0.088 & 1-0.045 2 3 4 5 6 7 8 0.135 -0.125 0EA: 20 KW x 200 mm - 4000K 45 = 0 D = 80 - 4000 X0.16 -0.114 m 0.93 D6 - 0.25 Ag =0 - 0.25A6 + 0.294 sg =-0.03 DE 2-0.117 m 0